# Underactuated Control of Multiple Soft Pneumatic Actuators via Stable Inversion

### v1

@@ Abstract
Soft grippers, with their inherent compliance and adaptability, show advantages for delicate and versatile manipulation tasks in robotics. This paper presents a novel approach to underactuated control of multiple soft actuators, specifically focusing on the synchronization of soft fingers within a soft gripper. Utilizing a single syringe pump as the actuation mechanism, we address the challenge of coordinating multiple degrees of freedom of a compliant system. The theoretical framework applies concepts from stable inversion theory, adapting them to the unique dynamics of the underactuated soft gripper. Through meticulous mechatronic system design and controller synthesis, we demonstrate both in simulation and experimentation the efficacy and applicability of our approach in achieving precise and synchronized manipulation tasks. Our findings not only contribute to the advancement of soft robot control but also offer practical insights into the design and control of underactuated systems for real-world applications.
@@
Underactuated control, algebraic control, soft actuator, soft robot, system perturbation.
@@ I Introduction
Soft robotics has become an emerging field, offering solutions for tasks that traditional robots struggle to accomplish. In particular, the inherent compliance and adaptability of soft actuators show advantages for applications requiring delicate manipulation [1, 2] and interaction with complex or unknown environments [3, 4]. Compared to rigid-bodied robotic hands, soft grippers stand out for their ability to conform to a wide range of object shapes and sizes, making them indispensable in domains such as medical robots [5] and human-robot interactions [6]. However, achieving precise control over the motion and coordination of multiple soft fingers within a gripper remains a challenge [7]. The particularly evident issue is underactuated control, where the number of actuators exceeds the number of control inputs, leading to restricted motions. Moreover, the air pump occupies space and has a certain weight, so reducing the number of inputs saves space and reduces weight and costs [8]. In this context, the development of effective underactuated control strategies tailored to the characteristics of soft robotics is essential to unlock their potential in real-world applications.
@@
Figure 1: The soft gripper has two fingers and is driven by a single syringe pump to achieve underactuated control via algebraic approaches. The control commands are generated in MATLAB/Simulink and are converted to PWM for the stepper motor in the syringe pump. The bending angles of both fingers are measured by the flex sensor embedded in each soft gripper.
@@
Despite the recent development of soft robot control, achieving precise control under the underactuated control framework for soft robots is seldom discussed and remains a challenge. A couple of works addressed the soft robot control issues by applying nonlinear controllers [9, 10], adaptive controllers [11, 12], and optimal controllers [7, 13]. Those control strategies enable high-performance control of soft robots with high degrees of freedom. However, the control of these systems becomes increasingly complex as the number of degrees of outputs exceeds that of inputs (underactuated control). Furthermore, soft materials have an uncertain nature and thus soft robots exhibit model uncertainty. When they work in unstructured environments, the robustness of the controller becomes more important. Addressing this challenge may require the development of novel control strategies that can effectively synchronize multiple soft actuators within a soft gripper, enabling precise and adaptive manipulation tasks in real-world scenarios.
@@
The primary objective of this paper is to develop control algorithms for achieving synchronization in multi-finger soft grippers. These mechatronic systems are designed and modeled to establish a non-square system matrix for single-input-multiple-output (SIMO) control scenarios. Our approach integrates both feedforward and feedback control loops. The feedforward control mechanism incorporates a stable model inversion technique that effectively synchronizes the motions of multiple soft fingers. Given the inherent uncertainty of soft materials, the feedback loop is adept at mitigating unexpected errors, noise, or disturbances that may arise between the mathematical model and the real system. Comprehensive validation of the control algorithms is conducted through simulations and experimentation. The theoretical framework underpinning these control algorithms is initially introduced in [14], where its efficacy is established by necessary and sufficient conditions. The contributions of this research lie in the innovative application of stable inversion control algorithms to address uncertain soft robotic systems. Notably, the proposed controller achieves synchronization of multi-finger soft grippers with a single input, thereby demonstrating the applicability of these control algorithms to SIMO control problems.
@@
To position our contributions, we compare our research with recent works. In our prior study [7], we employed individual optimal controllers for each finger within a multi-finger soft gripper (fully actuated) to attain precise and synchronized motions. In contrast, the present research adopts algebraic control algorithms alongside a single air pump to achieve synchronization across all fingers within the multi-finger soft gripper. Keppler et al. [15] established input coordination transformation that made the underactuated soft robotic systems become quasi-fully actuated systems. Although our feedforward control uses a similar concept, there is a feedback loop to cope with the uncertainty of the soft robotic systems that ensures robustness of the system. Pustina et al. [16] studied the controllability and stability of the underactuated soft robots. But our work focuses on the SIMO problem of synchronizing motions of multiple soft robots with a single input. Overall, this research studies the stable underactuation and synchronization of soft robots with robustness performance.
@@
The remainder of this paper is organized as follows. Section II introduces the mathematics preliminary and problem statement. Section III describes the full mechatronic design and algebraic controller design. Section IV evaluates the feasibility and applicability of the controller by simulations and experimentation. Section V discusses the experimental results and concludes the work.
@@ II-A Mathematics Preliminaries
The set $\mathbb{R}$ is a field of real number. Then, the set of polynomials in s with coefficients from $\mathbb{R}$ is denoted as $R(s)$. This set is a structured domain, a commutative ring over $\mathbb{R}$, labeled as $\mathbb{R}[s]$. The set, say $\mathcal{R}$, is an integral domain if this ring has an identity element and no zero divisors. A Euclidean Domain, labeled as $D$, is an integral domain paired with a Euclidean function denoted as $f:D\backslash\{0\}\rightarrow\mathbb{N}$, where $\mathbb{N}$ represents the set of integer numbers. To simplify, for any two elements $a$ and $b$ $\in D$ where $b$ is not equal to 0, there are two other elements $q$ and $r$ in $D$ such that $a=bq+r$, where $r$ can be 0 or the value of $f(r)<f(b)$.
Given that $n(s)$, $d(s)\in R[s]$ and $d(s)\neq 0$, the equation $n(s)/d(s)$ exists. The set of all these rational functions in $s$ over $R$ forms a field, denoted by $R(s)$ [17], and it holds that $R[s]\subset R(s)$. The sets of $n_{y}\times n_{u}$ matrices with elements in $R$, $R[s]$, and $R(s)$ are denoted by $R^{n_{y}\times n_{u}}$, $R[s]^{n_{y}\times n_{u}}$, and $R(s)^{n_{y}\times n_{u}}$ respectively. The R[s] is an Euclidean domain [14, 18]. In this paper, the order of a polynomial basis is defined as the degrees of its vectors.
@@
(Rouche-Capelli Theorem) Let $P\in\mathbb{R}(s)^{n_{y}\times n_{u}}$ be a matrix and $y\in\mathbb{R}(s)^{n_{y}\times 1}$ be a vector. If $rank(P)=rank([P~{}|~{}y])$, which is necessary and sufficient for $Pu=y$ to have a solution, the following statements hold:
C1. If $rank(P)=n_{u}\leq n_{y}$, the system has a unique solution.
C2. If $rank(P)=n_{y}<n_{u}$, the system has infinitely many solutions with $n_{u}-n_{y}$ free variables.
C3. If $rank(P)<min(n_{y},n_{u})$, the system has infinitely many solutions with $n_{u}-rank(P)$ free variables.
The definition in [19] is referenced to clarify the notation. The rank of a matrix over the field $\mathbb{R}(s)$ is defined as the maximum number of linearly independent subsets of its columns (or rows) with respect to $\mathbb{R}(s)$. This is denoted by $rank_{\mathbb{R}(s)}(P)$. If a set of vectors $v_{1},v_{2},...,v_{n_{u}}$ is linearly independent in the field $\mathbb{R}(s)$ if and only if the condition $\sum a_{i}v_{i}=0$ and $a_{i}=0$ for every $i$ with the scalars $a_{i}$ in $\mathbb{R}(s)$.
Suppose the rank of $P(s)\in\mathbb{R}(s)^{n_{y}\times n_{u}}$ is $r$ and $1\leq r\leq min(n_{u},n_{y})$. Let the columns of $P(s)=[p_{1},p_{2},...,p_{n_{u}}]$. Then, $\mathcal{L}(P)$ represents the set of $r$ linearly independent columns of $P(s)$ and defined as
$\begin{split}\mathcal{L}(P)=\{p_{i}\in P(s):1\leq i\leq r\}\end{split}$ (1)
with $p_{i}$ being linearly independent over $\mathbb{R}(s)$. We next introduce the subspace of $\mathbb{R}(s)^{n_{y}}$ spanned by these linearly independent columns
$\begin{split}\textnormal{Im}_{\mathbb{R}(s)}(P)=\biggl{\{}\sum_{i=1}^{r}{c_{i}}{p_{i}}:c_{i}\in\mathbb{R}(s),p_{i}\in\mathcal{L}(P)\biggl{\}}\subset\mathbb{R}(s)^{n_{y}}\end{split}$ (2)
The null space of $P(s)$ in the field $\mathbb{R}(s)$ is denoted as $\textnormal{Ker}_{\mathbb{R}(s)}(P)$. Let $\mathbb{R}[s]_{-}=\{n(s)\in\mathbb{R}[s]:\Re(\rho)<0,\forall{\rho}~{}s.t.~{}n(\rho)=0\}$ denote the set of polynomials in $\mathbb{R}[s]$ whose roots all have negative real parts and $\mathbb{R}(s)_{-}=\{n(s)/d(s)\in\mathbb{R}(s):d(s)\in\mathbb{R}[s]_{-}\}$ represents the set of rational functions in $\mathbb{R}(s)$ whose denominators belong to $\mathbb{R}_{-}$. The $\Re(\cdot)$ represents the real part of the given complex number and $\Im(\cdot)$ denotes the imaginary part of the number.
@@ II-B Problem Statement
The soft gripper is used as the plant for the proposed control algorithms. Although soft robots inherently possess nonlinear characteristics [20], the nonlinearity is often not obvious under restricted deformations. Thus, within the confines of limited deformations, soft robots can be effectively approximated using linear models. Consequently, the system can be described using a standard state-space representation [21].
Consider the multi-input-multi-output (MIMO) linear time-invariant (LTI) system
@@
where $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times n_{u}}$, $C\in\mathbb{R}^{n_{y}\times n}$, $x(t)$ and $x_{0}$ $\in\mathbb{R}^{n\times 1}$, $y(t)\in\mathbb{R}^{n_{y}\times 1}$, and $u(t)\in\mathbb{R}^{n_{u}\times 1}$. It is assumed that the (4) is Hurwitz and minimum phase system.
@@
The system given in (3) is minimal or is transformed into minimal realization.
The (3) can be solved mathematically and the derived equation in the Laplace domain ($\mathcal{L}\{f(t)\}=F(s)$ [22]) is shown below
The (4) can be expressed as
$\begin{bmatrix}Y_{1}(s)\\ \vdots\\ Y_{n}(s)\end{bmatrix}=\begin{bmatrix}P_{1}(s)\\ \vdots\\ P_{n}(s)\end{bmatrix}{U(s)}$ (5)
where $P_{i}(s)\in\mathbb{R}(s)^{n\times 1}$, $Y_{i}(s)\in\mathbb{R}(s)^{n\times 1}$, $i=1...n$, and $U(s)\in\mathbb{R}(s)$ so all elements in $P(s)$ and $Y(s)$ belongs to $\mathbb{R}(s)$. Also, the $Y(s)$ is in the range space of $P(s)$ and it is denoted as $Y(s)\in\textnormal{Im}_{\mathbb{R}(s)}(P)$. The system equation (5) can be analyzed in an algebraic manner. According to the Theorem 1, the (5) exists a unique solution and the $U(s)$ can be obtained. Since the system is Hurwitz and minimum phase, the solution, model inversion of $P(s)$, is stable. Thus, the solution exists and is stable.
@@ III Methodology
In this section, we thoroughly introduce the mechatronic system design including the soft pneumatic actuators [23, 21] and syringe pump design [24]. The research delves into the study of a soft gripper system formed by integrating those components. The soft pneumatic actuators serve as the fingers driven by the syringe pump and the algebraic control algorithms. The dynamical modeling of the systems based on mechanics and fluid dynamics theories is disclosed in Sec. III-B. The algebraic controller is designed based on the dynamical model as introduced in Sec. III-C.
Figure 2: The design of the mechatronic system can be seen in (a), (b), and (c), while the system modeling is observed in (d), (e), and (f). (a) visualizes how the optimal dimensional parameters are searched in a non-convex space. (b) illustrates the fabrication process of the soft pneumatic actuator and the flex sensor is embedded during the fabrication process. (c) shows the appearance of the syringe pump, and it is made of a commercial linear actuator and a commercial syringe. The (d) and (e) visualize how the structure of the soft actuator is approximated by a cantilever beam and how the bending angle is measured. The modeling schematic of the syringe pump is displayed in (g).

### v2

@@ Abstract
Soft grippers, with their inherent compliance and adaptability, show advantages for delicate and versatile manipulation tasks in robotics. This paper presents a novel approach to underactuated control of multiple soft actuators, explicitly focusing on the coordination of soft fingers within a soft gripper. Utilizing a single syringe pump as the actuation mechanism, we address the challenge of coordinating multiple degrees of freedom of a compliant system. The theoretical framework applies concepts from stable inversion theory, adapting them to the unique dynamics of the underactuated soft gripper. Through meticulous mechatronic system design and controller synthesis, we demonstrate the efficacy and applicability of our approach in achieving precise and coordinated manipulation tasks in simulation and experimentation. Our findings not only contribute to the advancement of soft robot control but also offer practical insights into the design and control of underactuated systems for real-world applications.
@@
Under-actuated control, stable inversion, soft actuator, soft robot, system perturbation.
@@ I Introduction
Soft robotics has become an emerging field, offering solutions for tasks that traditional robots struggle to accomplish. In particular, the inherent compliance and adaptability of soft actuators show advantages for applications requiring delicate manipulation [1, 2] and interaction with complex or unknown environments [3, 4]. Compared to rigid-bodied robotic hands, soft grippers stand out for their ability to conform to a wide range of object shapes and sizes, making them indispensable in domains such as medical robots [5] and human-robot interactions [6]. However, achieving precise control over the motion and coordination of multiple soft fingers within a gripper remains a challenge [7]. The soft fingers in this paper are pneumatically driven, and the air is supplied by air pumps. If the number of fingers exceeds the number of pumps, the system is under-actuated. Since the air pump is bulky, it is desired to minimize the number of air pumps [8, 9, 10].
@@
Figure 1: The soft gripper has two fingers and is driven by a single syringe pump to achieve underactuated control via model inversion. The control commands are generated in MATLAB®/Simulink and are converted to PWM for the stepper motor in the syringe pump. The bending angles of both fingers are measured by the flex sensor embedded in each soft gripper.
@@
Despite the recent development of soft robot control, achieving coordinated control under the underactuated control framework for soft robots is seldom discussed and remains a challenge. A couple of works addressed the soft robot control issues by applying nonlinear controllers [11, 12], adaptive controllers [13, 14], and optimal controllers [7, 15]. Those control strategies enable high-performance control of soft robots with high degrees of freedom. However, the control of these systems becomes increasingly complex as the number of degrees of outputs exceeds that of inputs (underactuated control) [8]. On the other hand, soft materials have an uncertain nature, and thus soft robots exhibit model uncertainty [16]. Various sources contribute to these uncertainties in modeling soft robots, which can be categorized into two types epistemic and statistical uncertainties [17]. Uncertainty can also arise from changes in the physical properties of soft materials. For example, polymer materials are susceptible to aging. Actuators or sensors made of polymer need time to fully cure and achieve chemical stability, during which their properties gradually shift [18]. Furthermore, prolonged use can alter the internal bonding structure of these materials [19]. Variations in the materials themselves and in manufacturing tolerances add another layer of uncertainty. When soft robots operate in unstructured environments [20], the robustness of the controller becomes more critical. Addressing this challenge may require the development of novel control strategies that can effectively coordinate multiple soft actuators within a soft gripper, enabling precise and adaptive manipulation tasks in real-world scenarios.
@@
The primary objective of this paper is to develop control algorithms for achieving coordination in multi-finger soft grippers, which are modeled as single-input-multi-output (SIMO) systems. Our approach integrates both feedforward and feedback control loops. The feedforward control mechanism incorporates a stable model inversion technique that effectively controls the motions of multiple soft fingers. Given the inherent uncertainty of soft materials, the feedback loop is adept at mitigating unexpected errors, noise, or disturbances that may arise between the mathematical model and the real system. Comprehensive validation of the control algorithms is conducted through simulations and experimentation. The theoretical framework underpinning these control algorithms is initially introduced in [21], where its efficacy is established by necessary and sufficient conditions. The contributions of this research lie in the innovative application of stable inversion control algorithms to address uncertain soft robotic systems. Notably, the proposed controller achieves coordination of multi-finger soft grippers with a single input, thereby demonstrating the applicability of these control algorithms to SIMO control problems.
@@
To position our contributions, we compare our research with recent works. In our prior study [7], we employed individual syringe pumps for each finger within a multi-finger soft gripper (full-drive) to attain precise and synchronized motions, so the number of air pumps is equal to the number of fingers. In contrast, the present research adopts stable model inversion alongside a single air pump to achieve coordination across all fingers within the multi-finger soft gripper. The algebraic-related method was proposed in [22], which established input coordination transformation that made the underactuated soft robotic systems become quasi-fully actuated systems. Although our feedforward control uses a similar concept, there is a feedback loop to cope with the soft robotic systems’ uncertainty that ensures the system’s robustness. Pustina et al. [23] studied the controllability and stability of the underactuated soft robots. But our work focuses on the SIMO problem of coordinating motions of multiple soft robots with a single input. Another major contribution is the experimental investigation and revelation of the parameter-varying nature of uncertainties in soft actuators. Furthermore, after conducting several experiments, a correlation between the uncertainty envelope and the bandwidth of underactuated control is established. Overall, this research studies the stable underactuation of soft robots with robust performance.
@@
The remainder of this paper is organized as follows. Section II introduces the mathematics preliminary and problem statement. Section III describes the full mechatronic design and proposed controller design. Section IV evaluates the feasibility and applicability of the controller by simulations and experimentation. Section V discusses the experimental results and concludes the work.
@@ II-A Mathematics Preliminaries
The set $\mathbb{R}$ is a real number field. Then, the set of all these rational functions in $s$ over $\mathbb{R}$ forms a field, denoted by $\mathbb{R}(s)$ [24]. The sets of $n_{y}\times n_{u}$ matrices with elements in $\mathbb{R}$ and $\mathbb{R}(s)$ are denoted by $\mathbb{R}^{n_{y}\times n_{u}}$ and $\mathbb{R}(s)^{n_{y}\times n_{u}}$ respectively.
The rank of a matrix $P$ over the field $\mathbb{R}(s)$ is defined as the maximum number of linearly independent subsets of its columns (or rows) [25]. This is denoted by $\textnormal{rank}_{\mathbb{R}(s)}(P)$. A set of vectors $v_{1},v_{2},...,v_{n_{u}}$ is linearly independent in the field $\mathbb{R}(s)$ if and only if the condition $\sum a_{i}v_{i}=0$ and $a_{i}=0$ for every $i$ with the scalars $a_{i}$ in $\mathbb{R}(s)$.
Suppose the rank of $P(s)\in\mathbb{R}(s)^{n_{y}\times n_{u}}$ is $r$ where $1\leq r\leq\text{min}(n_{u},n_{y})$. By choosing $P(s)$’s $r$ linearly independent columns, we can form a full-rank rational matrix as
[…]
with $p_{i}$ being linearly independent over $\mathbb{R}(s)$. Now, we can define the Image (Range) Set for the real-rational matrices in $s$ as follows.
$\begin{split}\textnormal{Im}_{\mathbb{R}(s)}(P)=\biggl{\{}\sum_{i=1}^{r}{c_{i}}{p_{i}}:c_{i}\in\mathbb{R}(s),p_{i}\in\mathcal{L}(P)\biggl{\}}\subseteq\mathbb{R}(s)^{n_{y}\times 1}\end{split}$ (2)
@@
(Rouche-Capelli Theorem) Consider $P\in\mathbb{R}(s)^{n_{y}\times n_{u}}$ with $\textnormal{rank}_{\mathbb{R}(s)}(P)=r$ and $Y\in\mathbb{R}(s)^{n_{y}\times 1}$. The solution(s) $U$ for the equation $PU=Y$ is exist if and only if
$\textnormal{rank}_{\mathbb{R}(s)}(P)=\textnormal{rank}_{\mathbb{R}(s)}\underbrace{([P(s)~{}:~{}Y(s)])}_{\in\mathbb{R}(s)^{n_{y}\times(n_{u}+1)}}=r$ (3)
Some notations used in the paper are: the $\Re(\cdot)$ represents the real part of the given complex number, and $\Im(\cdot)$ denotes the imaginary part of the number. $\mathcal{L}_{\infty}(j\mathbb{R})$ represents functions bounded on $\Re(s)$ = 0 including at $\infty$, and $\mathcal{RH}_{\infty}$ is the Hardy Space and denotes the set of $\mathcal{L}_{\infty}(j\mathbb{R})$ functions analytic in $\Re(s)>0$.
@@ II-B Problem Statement
The soft gripper, equipped with multiple fingers that can provide a human-like grasping experience, serves as the plant for the proposed control algorithms. Soft robotic systems, by their nature, exhibit nonlinear behaviors due to the compliance and deformability of their materials [16]. However, the nonlinearity of certain soft materials may not be evident under reasonably constrained deformations. Therefore, within the constraints of deformations, it is feasible to approximate the behavior of soft robots using linear uncertainty models [7, 26].
The multi-fingered soft gripper can be considered a type of multi-input-multi-output (MIMO) linear time-invariant (LTI) system where the state-space realization follows as
@@
where $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times n_{u}}$, $C\in\mathbb{R}^{n_{y}\times n}$, $x(t)$ $\in\mathbb{R}^{n\times 1}$, $y(t)\in\mathbb{R}^{n_{y}\times 1}$, and $u(t)\in\mathbb{R}^{n_{u}\times 1}$.
Equivalently, by using the Laplace transformation, (4) can be expressed with Transfer Function Matrices (TFM) as follows
@@
The system given in (5) is assumed to have the following properties:
  1. a.
    The system (5) is minimal, implying controllability and observability.
  2. b.
    The system (5) is the minimum phase and Hurwitz, meaning all poles and transmission zeros are in the left half of the complex plane.
  3. c.
    $Y(s)\in\textnormal{Im}_{\mathbb{R}(s)}(P)$, indicating that any output function is in the range space of $P$, making the output function achievable.
These assumptions ensure that the system is well-posed and stable, facilitating the design and analysis of the control strategies. These assumptions will be validated through analysis and assess their applicability under real-world conditions in Sec. IV-A.
To achieve a less complicated design, enhanced energy efficiency, and reduced weight, we can use a single input source to control all fingers with different dynamics and outputs. With this approach (having one input), the system described in (5) becomes a SIMO system
$\begin{bmatrix}Y_{1}(s)\\ \vdots\\ Y_{n_{y}}(s)\end{bmatrix}=\begin{bmatrix}P_{1}(s)\\ \vdots\\ P_{n_{y}}(s)\end{bmatrix}{U(s)}$ (6)
where $P_{i}(s)\in\mathbb{R}(s)$, $Y_{i}(s)\in\mathbb{R}(s)$, $i=1...n_{y}$, and $U(s)\in\mathbb{R}(s)$. Solving the algebraic equality for $U(s)$ gives us the exact left inverse. Since $P(s)$ (where $P(s)\neq 0)$ consists of a single-column real rational vector, we always have $\textnormal{rank}_{\mathbb{R}(s)}(P)=1$. Together with Assumption 1. c., the condition for the existence stated in Theorem 1 as
$\textnormal{rank}_{\mathbb{R}(s)}(P)=\textnormal{rank}_{\mathbb{R}(s)}{([P(s)~{}:~{}Y(s)])}=1$ (7)
is met, which ensures that (5) has a unique solution. On the other hand, since $P(s)$ has no unstable invariant zeros (Assumption 1 b.), the unique solution is also stable, i.e., U(s) $\in\mathcal{RH}_{\infty}$.
@@ III Methodology
In this section, we thoroughly introduce the mechatronic system design including the soft pneumatic actuators [27, 26] and syringe pump design [28]. The research delves into the study of a soft gripper system formed by integrating those components. The soft pneumatic actuators serve as the fingers driven by the syringe pump and the proposed control algorithms. The dynamical modeling of the systems based on mechanics and fluid dynamics is presented in Sec. III-B and the controller is designed in Sec. III-C.
Figure 2: The design of the mechatronic system can be seen in (a), (b), and (c), while the system modeling is observed in (d), (e), and (f). (a) visualizes how the optimal dimensional parameters are searched in a non-convex space. (b) illustrates the fabrication process of the soft pneumatic actuator and the flex sensor is embedded during the fabrication process. (c) shows the appearance of the syringe pump, and it is made of a commercial linear actuator and a commercial syringe. The (d) and (e) visualize how the structure of the soft actuator is approximated by a cantilever beam and how the bending angle is measured. The modeling schematic of the syringe pump is displayed in (f).